# Visualizationary: Automating Design Feedback for Visualization Designers using LLMs

### v1

@@ 5.4 Data Collection
P11 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
P3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
P9 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Examples of visualization design evolution. These images represent the five most significant versions (as identified by the participants) during the design process of creating a new visualization artifact from scratch. We include data for one representative user from each of the novice (P11), intermediate (P3), and expert (P9) groups. Complete data for all participants can be found in the supplemental material.
@@ 6.1 Design Evolution
Fig. 4 presents visualization snapshots from three representative participants: novice, intermediate, and expert groups. We only show three examples due to page limits; the complete snapshots and descriptions from all participants can be found in our supplementary material. We asked participants to explain the changes created in each iterative step during the interview and describe their intent in making changes below. Below we discuss the design evolution for each participant as follows: the visualization, authoring tool, and dataset used, a summary of each version, including the feedback from Visualizationary, the changes made, and—if given—their rationale.
@@
Participant 11 (novice). P11 created a stacked area chart (on Microsoft Excel’s recommendation) to visualize daily activities.
@@
  • Version 1$\rightarrow$2: Feedback: Visualizationary recommended bar charts to enhance value comparisons. Revision: P11 changed from stacked line chart to a stacked bar chart.
@@
  • Version 2$\rightarrow$3: Feedback: Visualizationary noted that there is too much focus of salience towards the center. Revision: P11 adjusted the chart’s size ratio and incorporated a new line chart, which displayed the cumulative average hours spent on a specific activity (“work”). P11 introduiced a vertical axis to represent the line value in the chart. Rationale: P11 wanted to examine influences on work.
@@
  • Version 3$\rightarrow$4: Feedback: Again, Visualizationary observed excessive concentration of salience in the central area. Revision: P11 re-adjusted the chart dimensions, and highlighted legend and title.
@@
  • Version 4$\rightarrow$5: Feedback: Saliency is overly focused on the center. Revision: P11 increased title and legend font sizes.
@@
Participant 3 (intermediate). P3 wanted to analyze their monthly spending based on two factors: (1) category, and (2) payment method. They used Tableau to author the visualization.
@@
  • Version 1$\rightarrow$2: Feedback: Visualizationary asserted that there is too much focus on text. Revision: P3 changed the order of the hierarchy to remove repetitive labels.
@@
  • Version 2$\rightarrow$3: Feedback: Again, Visualizationary stated that there is too much focus on text. Revision: P3 removed the “payment method” category, consolidated its values into a stacked bar chart, and labeled each bar segment with its corresponding payment method.
@@
  • Version 3$\rightarrow$4: Feedback: And once again, Visualizationary said that there is still too much focus on text. Revision: P3 eliminated the text alongside the bar chart and introduced a legend instead.
@@
  • Version 4$\rightarrow$5: Feedback: And again, Visualizationary maintained that there is still too much text. Revision: P3 rotated the bars vertically and separated the labels for categories and subcategories.
@@
Participant 9 (expert). Using Tableau, P9 chose to display a multinational company’s profits in Asia on a geographic map.
@@
  • Version 2$\rightarrow$3: Feedback: The salience map in Visualizationary directed attention country names, but several were unrelated to the data. Revision: Country names were added on top of each circle, and irrelevant country names originally on the map were removed.
@@
  • Version 3$\rightarrow$4: Feedback: Visualizationary noted that textual content would make the chart more readable. Revision: P9 relocated country names to circle centers, added explanatory boxes, and differentiated ocean colors for clarity.
@@
  • Version 4$\rightarrow$5: Feedback: The visual entropy filter in the tool did not detect some circles. Revision: Strokes were added to circles.
Table II: Questions for post-study interviews. We asked participants a total of 7 questions about their experience in using Visualizationary to understand the role of Visualizationary in the design process, to unveil their decision-making process in acting on feedback, and to evaluate the tool.
Question Type Subject Questions Role Analyze How was your chart problem detection aided by Visualizationary? Clarify How did Visualizationary assist in clarifying chart guidance or analysis? Guide How was the guidance from Visualizationary for visualization design? Track How did tracking influence your design modifications? Decision Judgment In general, when do you act on feedback or decline it? Evaluation System What are the pros/cons of Visualizationary when the LLM bases answers on established visualization design knowledge? Impression Overall, how did you like the tool?
@@ 6.2 Post-Study Interview
After the experiment, we conducted a post-interview session to examine closely the iterative improvements of the design and the role of Visualizationary in making the changes. Our initial set of questions (Table II) aimed to reacquaint the participants with the changes their designs underwent and to understand the underlying reasons behind them. For every design refinement, we asked participants to elaborate on the following: (1) the changes in the design, (2) the reasons for the changes in the design, and (3) the role of Visualizationary in the changes. We then transitioned to understand their experiences in using Visualizationary during the design process, and their perspective on design feedback mechanisms. Our goals were the following:
  1. 1.
    Role. Understand Visualizationary’s role in ACGT.
  2. 2.
    Decision. Identify when designers accept/reject feedback.
  3. 3.
    Evaluation. Evaluate strengths and weaknesses.
Detailed below are the questions presented to our participants.
Table III: Evaluation of design evolution. We asked three external experts to evaluate the improvements in the designs generated by 13 participants. Scores are measured using a 1–5 Likert scale, where 1 represents a significant decline, 3 is neutral, and 5 is a significant improvement. Each value shows the mean value of scores from external evaluators; values in parentheses are standard deviations.
Part. P1 P2 P3 P4 P5 P6 P7 P8 P9 P10 P11 P12 P13 Avg Avg. 4.67 2.67 3.67 3.00 5.00 4.33 2.67 3.00 4.67 4.00 3.67 2.67 4.00 3.69 S.D. (0.58) (0.58) (0.58) (0.00) (0.00) (0.58) (0.58) (1.73) (0.58) (0.00) (0.58) (0.58) (0.00) (0.83)
@@
Role for Analyze: Expert Analyst. Visualizationary is capable of not only detecting issues, but also in providing relevant knowledge that only experts can have. Some of the observed comments were aspects participants had not initially considered, yet were contributing to the improvement of the visualization. In other instances, the tool’s feedback was on aspects participants were familiar with or had previously encountered, but which were typically overlooked during visualization design, such as the color deficiency filters. P2 said, “[the tool] is capable of showing problems that I was not aware of, although I should have been.” Upon diagnosis, the tool supplies relevant data from visualization design knowledge, which, as P8 noted, is information that even experts might occasionally miss. Said P8, “even experts may not be aware of those specific details about visualization principles.” However, there were also times when the tool misdiagnosed the chart image, resulting in potentially out-of-context feedback.
Role for Clarify: Organizer. Because the output of this report is generated by ChatGPT, it is predominantly textual. The hierarchical layout of the design allowed participants to view the summary of each section, selectively pick the issues that they found relevant and made sense, and drill down to the details they were interested in. 10 out 13 participants felt that this capability helped them be more organized.
Role for Guide: Guidepost. Given a detectable issue, Visualizationary is capable of providing general suggestions on how to address them, but does not always deliver specific actionable feedback. This means that human intervention is still required to find a solution. Participants also noted variability in the guidance: some feedback was easy to implement whereas other feedback required a substantial amount of effort to address. Easily addressed were tasks that can easily be tested, such as editing colors, adding texts, and so on. All participants attempted to address issues raised by the tool. However, a few novice and intermediate participants found it difficult to address more complex problems, such as diverting or attracting attention in the saliency map. For example, P3 spent all their iterations to balance the salience of a chart, and P7 wanted the salience to focus on a particular area.
@@
Role for Track: Logbook. The tracking capabilities play as a role similar to a design “logbook.” Tracking changes per design iteration allowed some participants to easily catch up to where the design was left off. A notable observation was that it also functioned as an evaluation tool that measures whether there were improvements between the current and the past version. P3, whose main goal throughout the study was to alleviate salience in textual areas, said, “I consistently checked the rate of salience in the tracking interface. The drop in numbers reassured me that I was on the right path. The positive feedback felt like a reward for the adjustments I had made.”
@@
On the other hand, there were also a few participants who did not use tracking. They uploaded the designs one after another without much delay. P2 was an example, noting “Because I remember what needs to be changed, I did not look into the tool.”
Criteria for Accepting Feedback. As one might expect, there were no participants that accepted all the feedback given by Visualizationary for their designs. Designers clearly cannot make changes based on what they cannot understand (or disagree with). For example, P7 did not understand how to create a bar chart using the dataset that they had. Second, the designer must also agree to some extent that the feedback would enhance the design. For example, some designers disregarded feedback on color vision deficiency when the visualization was merely intended for their own use. When the designer was unsure of whether there would be improvements, if the feedback was easily actionable, then they just tried it and took the feedback when it did.
@@
General Assessment. Participants considered the tool as an assistant that provided expert-driven feedback. Many people originally thought that it would be difficult to provide insights, but were surprised at its capability to help in visualization design. Furthermore, the use of ChatGPT created a sense of a “human” responding to the design, which made some participants less isolated during the design process.
However, some areas could be enhanced. One main comment was that despite the hierarchical presentation, several participants felt somewhat overwhelmed by the volume of text in the report. Some participants felt that with the help of visual images and text combined, this tool would generate more effective feedback.
6.3 Design Evaluation
We here present the results of the design assessment by external reviewers, which is the final step of our study procedure (see Sec. 5.3). We recruited three senior visualization researchers who all have more than 7 years of experience to evaluate the designs conducted by the participants. Table III shows the overall ratings on the improvement of visualization designs conducted by our participants. Note that these values are not on the quality of the design but on the improvement of the design. On the whole, evaluators largely believed that the participants’ designs improved from their starting point, averaging a score of 3.69 (3 is neutral). P2, P7, and P12’s designs were modestly below the neutral mark. Evaluators observed consistency in P2’s designs, with a slight decrease in color quality. For P7, while adding a title was beneficial, the chart’s readability declined. P12 enlarged the chart for improved readability, but received mixed reviews about the choice of color. We also compare the improvements per different expert levels. We find that the mean score of improvements from the novice group is $3.34$, from the intermediate group is $3.92$, and from the expert group is $4.11$. While we observe positive signs from all groups, we see more improvements from intermediate and expert groups.

### v3

@@ 5.4 Data Collection
N1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
N2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
N3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
N4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
N5 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
N6 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Examples of visualization design evolution (Novices). These images represent the five most significant versions (as identified by the participants) of novice visualization designers during the design process of creating a new visualization artifact from scratch. The explanations of these design iterations can be found in the supplemental material.
Version 1 Version 2 Version 3 Version 4 Version 5
I1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
I2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
I3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
I4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
E1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
E2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
E3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Examples of visualization design evolution (Intermediate and expert designers). These images represent the five most significant versions (as identified by the participants) of intermediate and expert visualization designers during the design process of creating a new visualization artifact from scratch. The explanations of these design iterations can be found in the supplemental material.
@@ 6.1 Design Evolution
[…]
ID Iteration 1 Iteration 2 Iteration 3 Iteration 4
[…]
Here, we describe how the participants’ designs evolved. First, we provide an overview of their experimental setup. Table II details the tools used, dataset themes, chart types, and the time spent iterating each design. We found that a variety of tools (6 in total) are employed regardless of expertise level, and the datasets covered diverse topics. This justifies our decision to upload visualizations as image files, since this approach keeps the feedback independent of the tool. Bar charts were the most common chart type, adopted by eight participants, though nine different visualization types appeared in total. The time spent is calculated as the interval between the first and last upload. While we cannot determine exactly how much effort was allocated to each design within this period, a few patterns emerged. Participants worked in one long session, and it lasted approximately 90 to 150 minutes.
@@
Second, we present information about participants’ design evolution processes. Figs. 3 and 4 show visualization snapshots from all 13 participants, categorized into novice, intermediate, and expert groups. Larger snapshots and full descriptions appear in the supplementary material. During interviews, we asked participants to explain the changes they made in each iterative step and to discuss their intentions. Below we discuss the design evolution for three selected participants, N1, I2, and E1 as follows: the visualization, authoring tool, and dataset used, a summary of each version, including the feedback from Visualizationary, the changes made, and—if given—their rationale. Results for participants not covered here can be found in the supplementary material.
@@
Participant N1 (novice). N1 created a stacked area chart (on Microsoft Excel’s recommendation) to visualize daily activities.
@@
  • Version 1$\rightarrow$2: Feedback: Visualizationary recommended bar charts to enhance value comparisons. Revision: N1 changed from stacked line chart to a stacked bar chart.
@@
  • Version 2$\rightarrow$3: Feedback: Visualizationary noted that there is too much focus of salience towards the center. Revision: N1 adjusted the chart’s size ratio and incorporated a new line chart, which displayed the cumulative average hours spent on a specific activity (“work”). N1 introduced a vertical axis to represent the line value for examining influences on work.
@@
  • Version 3$\rightarrow$4: Feedback: Again, Visualizationary observed excessive concentration of salience in the central area. Revision: N1 re-adjusted chart dimensions and highlighted labels.
@@
  • Version 4$\rightarrow$5: Feedback: Saliency is focused in unimportant areas. Revision: N1 increased title and legend font sizes.
@@
Participant I2 (intermediate). I2 wanted to analyze their monthly spending based on two factors: (1) category, and (2) payment method. They used Tableau to author the visualization.
@@
  • Version 1$\rightarrow$2: Feedback: Visualizationary asserted that there is too much focus on text. Revision: I2 changed the order of the hierarchy to remove repetitive labels.
@@
  • Version 2$\rightarrow$3: Feedback: Again, Visualizationary stated that there is too much focus on text. Revision: I2 removed the “payment method” category, consolidated its values into a stacked bar chart, and labeled each bar segment with its corresponding payment method.
@@
  • Version 3$\rightarrow$4: Feedback: And once again, Visualizationary said that there is still too much focus on text. Revision: I2 eliminated the text in the bar chart and introduced a legend.
@@
  • Version 4$\rightarrow$5: Feedback: And again, Visualizationary maintained that there is still too much text. Revision: I2 rotated the bars vertically and separated the category labels.
@@
Participant E1 (expert). Using Tableau, E1 chose to display a multinational company’s profits in Asia on a geographic map.
@@
  • Version 2$\rightarrow$3: Feedback: The salience map in Visualizationary directed attention to country names, but several were unrelated to the data. Revision: Country names were added to each circle, and irrelevant country names were removed.
@@
  • Version 3$\rightarrow$4: Feedback: Visualizationary noted that textual content would make the chart more readable. Revision: E1 relocated country names to circle centers, added explanatory boxes, and differentiated ocean colors for clarity.
@@
  • Version 4$\rightarrow$5: Feedback: Scanner Deeply did not detect some circles. Revision: Strokes were added to circles.
6.2 Feedback Usage Patterns
Below we present a summary of feedback usage patterns. Table III outlines the filters each participant used during the design process. To begin with, we describe iteration patterns, and then we also analyze cases when participants do not take the feedback.
Iteration patterns. Across the 13 participants, each design iteration typically incorporated one or two feedback changes, occasionally reaching three in a single iteration. Of the 68 total feedback instances, 33 were related to “Scanner Deeply,” with 18 of these examining whether salience aligned with the designer’s intended focal points. Notably, salience-related feedback often required multiple iterations to complete, possibly because shifting salience in a single step is not straightforward. Chart recommendations appeared only five times—primarily in the first iteration. This seems to reflect the tendency to make major design choices (e.g., chart type) prior to making minor design choices [58]. The chartjunk filter was never triggered, as no chartjunk elements were introduced in the participants’ designs. Although usage patterns were relatively consistent across skill levels, a slight decrease in implemented feedback was observed from early to late iterations (an average of 1.46 changes in the first iteration versus 1.23 in the last), which may indicate that designs converged over time, reducing the need for additional modifications.
When participants do not listen to the feedback. We observe several instances where participants noted the system’s feedback but ultimately chose not to implement it. The most frequent reason was that the model did not fully align with the user’s intent. When I4 was advised to use additional colors for separate categories, they declined because only two colors were necessary for their comparative task. In other cases, users felt the system’s suggestions conflicted with their design objectives. For instance, N4 wanted to include a grid for comparison purposes, despite realizing it might reduce salience in key areas. Some participants also proceeded with their own experiments unrelated to any feedback, such as N1 adding a line chart to meet personal objectives. Finally, another example is when the system presents incorrect recommendations. For instance, participant I1 was advised to replace a heatmap with a bar chart but did not know how to create this chart for their data.
Overall, these observations indicate that the system is not as effective when users have their own objectives, as it currently lacks the capacity to capture their intentions. They also reveal that some filters within the system are not optimal.
Table IV: Expert evaluation of visualization designs. We deployed 3 seasoned visualization experts to evaluate visualization designs authored by 13 participants. The scores from second to fifth columns show the evaluation of visualization designer per iteration. We present the trends of scores per participant in the sixth column. We also asked the experts to pick the best design from each designer’s iterations. This is shown in the seventh column. Finally, the last column shows the overall improvement scores of visualization designs. The three numbers in brackets in the third and fifth column represent the scores obtained by the experts, sorted per different expert evaluator. The 1–5 Likert scale is defined as follows: 1 = significant decline in quality, 3 = neutral, and 5 = significant improvement in quality.
ID Expert evaluation scores per iteration Trends Votes for the best version Overall score (V1 - V5)
V1 $\rightarrow$ V2 V2 $\rightarrow$ V3 V3 $\rightarrow$ V4 V4 $\rightarrow$ V5
N1    3.00 (3/4/2)    3.00 (5/2/4)    3.67 (3/4/4)    3.33 (3/3/4)
{sparkline}
[…]
1 1 1
3.67 (4/3/4)
N2    2.33 (4/1/2)    2.67 (3/2/3)    3.00 (4/3/2)    3.00 (4/3/2)
{sparkline}
[…]
2 1
2.67 (3/3/2)
N3    3.33 (3/3/4)    1.67 (1/2/2)    4.67 (4/5/5)    3.33 (4/2/4)
{sparkline}
[…]
1 2
4.00 (4/4/4)
N4    3.00 (3/4/2)    2.67 (2/3/3)    2.67 (3/3/2)    3.00 (4/3/2)
{sparkline}
[…]
1 1 1
2.67 (3/3/2)
N5    4.33 (4/4/5)    2.33 (3/2/2)    3.00 (4/1/4)    3.00 (2/3/4)
{sparkline}
[…]
1 2
3.00 (3/3/3)
N6    4.33 (4/4/5)    4.33 (4/5/4)    4.00 (3/5/4)    4.00 (4/4/4)
{sparkline}
[…]
3
4.00 (4/4/4)
I1    3.67 (3/4/4)    3.00 (4/3/2)    2.67 (2/3/3)    3.33 (3/4/3)
{sparkline}
[…]
1 1 1
2.67 (3/3/2)
I2    4.33 (4/4/5)    4.00 (5/3/4)    3.67 (2/4/5)    2.33 (2/3/2)
{sparkline}
[…]
3
3.67 (4/4/3)
I3    3.33 (4/2/4)    3.33 (4/3/3)    2.33 (1/3/3)    5.00 (5/5/5)
{sparkline}
[…]
3
5.00 (5/5/5)
I4    4.00 (4/5/3)    4.00 (4/4/4)    4.33 (5/4/4)    4.33 (3/3/4)
{sparkline}
[…]
2 1
4.33 (5/4/4)
E1    3.00 (2/3/4)    4.33 (4/5/4)    4.00 (5/3/4)    3.67 (3/4/4)
{sparkline}
[…]
3
4.67 (4/5/5)
E2    3.00 (2/5/2)    3.33 (3/3/4)    4.33 (5/4/4)    4.33 (4/4/5)
{sparkline}
[…]
3
3.00 (1/4/4)
E3    4.33 (4/4/5)    3.67 (3/4/4)    4.67 (5/5/4)    3.00 (3/3/3)
{sparkline}
[…]
2 1
4.67 (4/5/5)
Table V: Feedback evaluation per filter types. We present the average feedback score for each filter type, based on iteration-level scoring. We also provide a summary of the effectiveness and limitations of the filters used in Visualizationary per topic, based on the comments provided by the expert evaluators. The numbers in brackets indicate the corresponding count of feedback instances for each score. Overall, designers improved when using the feedback, although novices found the CVD-related feedback more challenging to apply.
Novices Intermediates Experts Effectiveness Limitations
[…]
6.3 Design Evaluation
We here present the results of the design assessment by external reviewers, which is the final step of our study procedure (see § 5.3). We recruited three senior visualization researchers who all have more than 7 years of experience to evaluate the designs conducted by the participants. Table IV shows various evaluations on the improvement of visualization designs conducted by our participants. Columns from second to fifth show the evaluation of visualization designer per iteration. The sixth column shows the increasing/decreasing trend of the scores per participant. The seventh column shows which design the experts judged to be the best among each designer’s iterations. Finally, the last column shows the overall improvement scores of visualization designs. Note that the values in the second to fifth and the last columns are not on the quality of the design but on the improvement of the design. Also, we analyze the evaluation scores, effectiveness, and limitations per feedback types. This is shown in Table V.
On the whole, evaluators largely believed that the participants’ designs improved from their starting point, averaging a score of 3.69 (3 is neutral). N2, N4, and I1’s designs were modestly below the neutral mark. Evaluators observed consistency in N4’s designs, with a slight decrease in color quality. For I1, while adding a title was beneficial, the chart’s readability declined. N2 enlarged the chart for improved readability, but received mixed reviews about the choice of color. We also compare the improvements per different expert levels. We find that the mean score of improvements from the novice group is $3.34$, from the intermediate group is $3.92$, and from the expert group is $4.11$. While we observe positive signs from all groups, we see more improvements from intermediate and expert groups. Below, we provide detailed analyses of the tables.
When feedback is effective. We identified several ways in which the system’s feedback proved effective. First, as designers gained more expertise, they used the salience feature more effectively, which led to an improved data ink ratio. This was also noted by the evaluators. Most issues raised by the system were resolved within one or two iterations, although salience maps sometimes required additional effort. Even novice users managed salience related challenges by comparing designs across iterations and using the tracker interface (Fig. 2 (D)) together with archives for more detailed heatmap comparisons. In addition, feedback on text placement and color usage led to tangible improvements.
These findings indicate that our ACG feedback framework, as well as feedback in textual form, was effective across various skill levels. We describe the role of the tracker interface (T) in § 6.4.
When feedback is not effective. Although the system demonstrated effectiveness in many situations, it was not universally successful. One reason is that novices did not receive feedback that was sufficiently direct or actionable. In practice, novices had difficulty adapting color palettes for color vision deficiencies, which indicates that selecting and applying a limited set of colors can be challenging for less experienced designers. Some also struggled with salience-related tasks, such as precisely shifting attention to specific chart elements. In the future, we plan to explore more accessible ways to recommend color palettes for CVD and to provide more direct salience-related feedback.
Furthermore, the system focuses on general communicative purposes, and thus does not yet account for specialized design intentions. For example, I1 prioritized data analysis rather than optimizing the chart for communicative goals. Although the expert evaluators were unaware of this, it may still have been appropriate for the designer’s specific objectives.
Moreover, we observe cases where the filters in the system occasionally malfunction. For example, salience may appear concentrated in a region with no chart elements (N6), or, if the chart is too complex (I1), Deplot might fail to read the data correctly and recommend an inappropriate chart. Although such cases are relatively not common, they reduce confidence in the filters.
Best version from iterations. To determine which iteration yielded the most polished design, we asked evaluators to select the best version out of five. Our overview shows that when a design’s score is lower, evaluator opinions are more divided, whereas a higher score reflects consensus favoring the fifth version. However, the final design is not always the best for several reasons. First, if feedback-driven changes do not result in noticeable improvements, evaluators may choose different versions. For instance, in participant I1’s case, the design changes were minor and barely discernible, leading three evaluators to select three different versions. N4 exhibited a similar pattern. Second, designers’ goals can sometimes introduce unintended changes that degrade the overall design from an outside perspective. For example, near the end of participant E3’s process, an element became overly highlighted; evaluators, unaware of this intention, did not favor that final iteration. Third, when changes are trivial, observers may not detect a meaningful difference between one iteration and the next. For example, in participant N3’s final iteration, only a single red grid line was added to enhance clarity, which did not substantially shift evaluators’ perceptions.
Overall, when design changes are clear and demonstrably beneficial, evaluators’ judgments tend to converge. Conversely, if the changes are less distinct or their benefits uncertain, their evaluations diverge. We also observe that some designers have their own goals, and evaluators assessed them as a communicative visualization without knowing the designers’ intents.
Variations in evaluators’ scores. In some cases, the evaluators’ scores converged, yet there were also instances of considerable divergence. Overall, the variance in final scores was generally minimal: most evaluators agreed or differed only slightly. However, participant E2 presented an exception, where two evaluators gave a score of 4 and one assigned a 1. The evaluator who gave a 1 stated, “I do not see how changing from a heatmap to a bar chart is an improvement. I might even prefer the heatmap.” This discrepancy appears related to the communicative aspect of visualization, in which the evaluator’s limited understanding of the designer’s intentions played a role.
We observed even greater variations in the evaluators’ assessments across different iterations. Despite having solid expertise in visualization, each evaluator brought unique preferences and standards for communicative visualization. For example, moving from V2 to V3 in I2’s design, one evaluator was neutral about introducing multiple colors (assigning a 3), another greatly approved (giving a 5), and a third specifically praised the altered chart type (awarding a 4). Similar disagreements arose when some evaluators judged the overall positive effect of a design change, whereas others compared only the previous and current versions. One evaluator noted, “ To judge whether a change is good or bad, we need to consider the entire sequence of modifications. Design is not always a greedy process. Can we dismiss a change as having low value just because it looks unfavorable in one instance?
These findings underscore the complexity of evaluation when preferences vary. They also highlight that evaluators, unfamiliar with the designer’s intention, may reach divergent conclusions.
@@ 6.4 Post-Study Interview
After the experiment, we conducted a post-interview session to examine closely the iterative improvements of the design and the role of Visualizationary in making the changes. For every design refinement, we asked participants to mainly elaborate on the role of Visualizationary in the changes. We then transitioned to understand their experiences in using Visualizationary during the design process, and their perspective on design feedback mechanisms. Detailed below are the summarized comments.
@@
Role of the feedback. Visualizationary not only detects issues but also provides expert-level knowledge that many participants had not initially considered. In some cases, the system highlighted overlooked concerns, such as color deficiency filters. N4 noted, “[the system] is capable of showing problems that I was not aware of, although I should have been,” while E2 remarked, “even experts may not be aware of those specific details about visualization principles.” However, the system occasionally misdiagnosed the chart image, leading to feedback that was out of context.
Moreover, Visualizationary could provide a more detailed, actionable solutions, as some complex tasks still required deep human decisions. Although all participants attempted to address the system’s feedback, some novice- and intermediate-level designers found it difficult to implement suggestions that were not specific enough, such as redistributing saliency attention. For instance, I2 devoted every iteration to balancing a chart’s salience, and I1 sought to focus salience on a particular area, underscoring the varying difficulty of applying automated feedback.
@@
Role of tracker. The consistent use of the tracker interface (T) for iterative refinements, especially when a problem could not be resolved in a single step, further underscores its supportive role in the design process. We find that most participants that use virtual eyetracker have the experience of using the tracker interface to track their improvements (11/13). For people that use saliency maps, it functioned as a comparison system that measures whether there were improvements between the current and the past version. I2, whose main goal throughout the study was to alleviate salience in textual areas, said, “I consistently checked the rate of salience in the tracking interface. The drop in numbers reassured me that I was on the right path. The positive feedback felt like a reward for the adjustments I had made.” Furthermore, tracking changes per iteration allowed participants to catch up (3/13).
@@
There were also a few participants who did not use tracking. They uploaded the designs one after another without much delay. N4 was an example, noting “Because I remember what needs to be changed, I did not look into the system.”
Role of the textual report. Because the output of this report is generated by an LLM, it is predominantly textual. The hierarchical layout of the design allowed participants to view the summary of each section, selectively pick the issues that they found relevant, and drill down to the important details. 10 out 13 participants felt that this capability helped them be more organized, though some still felt overwhelmed at the volume of text.
@@
General assessment. Participants considered the system an assistant providing expert feedback. Many people originally thought that it would be difficult to provide useful insights, but were surprised at its capability to help the design. Furthermore, the use of an LLM created a sense of a “human” response, which made some participants feel less isolated during the design process.
However, some areas could be enhanced. One key observation was that, despite the hierarchical structure, participants still found the text volume overwhelming.